White Tiger Electronics produces CD players using an automated assembly line process. The standard cost of CD players is

​$156

per unit​ (labor,

​$33

​materials,

​$69 and​ overhead,

​$54

The sales price is

​$300

per unit.

A. To achieve a 14 percent multifactor productivity improvement by reducing materials costs​ only, one needs to reduce the materials cost by _____ percent.

B. To achieve a 14 percent multifactor productivity improvement by reducing labor costs​ only, one needs to reduce the labor cost by ________

percent.

C. To achieve a 14 percent multifactor productivity improvement by reducing labor costs​ only, one needs to reduce the overhead costs by ________

percent.
Answer #1

Here,

Given Data, Value of Output = USD 300,

Value of Input:

Labor Input: USD 33,

Material Input: USD 69

Overhead Input: USD 54

Total Value of Input = Labor Input + Material Input + Overhead Input = 33 + 69 + 54 = USD 156

Multifactor Productivity = Output/ Total Input = USD 300/USD 156 = 1.923      

14% Multifactor productivity improvement = 1.14 * 1.923 = 2.192,

Now given that Multifactor productivity = 2.192, value of output = USD 300,

So Value of Total input = Output/Multifactor productivity = USD 300/2.192 = 136.86,

Thus, the cost of Inputs must decrease by USD 156 - USD 136.86 = USD 19.14

A. To achieve a 14 percent multifactor productivity improvement by reducing materials cost only, one needs to reduce the materials cost by USD 19.14/USD 69 = 27.74 percent.

B. To achieve a 14 percent multifactor productivity improvement by reducing labor costs​ only, one needs to reduce the labor cost by USD 19.14/ USD 33 = 58 percent.

C. To achieve a 14 percent multifactor productivity improvement by reducing overhead cost only, one needs to reduce the overhead costs by USD 19.14/USD 54 = 35.44 percent.